a squared = (x+3)
b squared = 4(x+2)
c squared = 625
I have got to:
xsq+6x+9+4xsq+16x+16=625
then get stuck

Answer by stanbon(50455)

a squared = (x+3)
b squared = 4(x+2)
c squared = 625
----
Pythagoras:
a^2 + b^2 = c^2
x+3+4x+8 = 625
5x+11 = 625
5x = 614
x = 122.8
==============
Cheers,
Stan H.
==============